### v1

# Decoupled structure-preserving discretization of incompressible MHD equations with general boundary conditions

Abstract
In the framework of a mixed finite element method, a structure-preserving formulation for incompressible MHD equations with general boundary conditions is proposed. A leapfrog-type temporal scheme fully decouples the fluid part from the Maxwell part by means of staggered discrete time sequences and, in doing so, partially linearizes the system. Conservation and dissipation properties of the formulation before and after the decoupling are analyzed. We demonstrate optimal spatial and second-order temporal error convergence and conservation and dissipation properties of the proposed method using manufactured solutions, and apply it to the benchmark Orszag-Tang and lid-driven cavity test cases.
keywords:
MHD , mixed finite element method , structure-preserving discretization , general boundary conditions
journal: arXiv
[…]
[1]organization=School of Mathematics and Computing Science, Guilin University of Electronic Technology, city=Guilin, country=China
[…]
[2]organization=Delft Institute of Applied Mathematics, Delft University of Technology, city=Delft, country=the Netherlands
[…]
[3]organization=ICA, Université de Toulouse, ISAE–SUPAERO, INSA, CNRS, MINES ALBI, UPS, city=Toulouse, country=France
[…]
[4]organization=Faculty of Aerospace Engineering, Delft University of Technology, city=Delft, country=the Netherlands
1 Introduction
Given a bounded, contractible domain $\Omega\subset\mathbb{R}^{d}$, $d\in\left\{2,3\right\}$ with a Lipschitz boundary $\partial\Omega$, the incompressible, constant density magnetohydrodynamic (or simply incompressible MHD) equations [1] written as
(1a) $\rho\left[\partial_{t}\bm{u}+\left(\bm{u}\cdot\nabla\right)\bm{u}\right]-\tilde{\mu}\Delta\bm{u}-\bm{j}\times\bm{B}+\nabla p$ $=\rho\bm{f},$
(1b) $\nabla\cdot\bm{u}$ $=0,$
(1c) $\partial_{t}\bm{B}+\nabla\times\bm{E}$ $=\bm{0},$
(1d) $\bm{j}-\sigma\left(\bm{E}+\bm{u}\times\bm{B}\right)$ $=\bm{0},$
(1e) $\bm{j}-\nabla\times\bm{H}$ $=\bm{0},$
(1f) $\bm{B}$ $=\mu\bm{H},$
govern the dynamics of fluid velocity $\bm{u}$, electric current density $\bm{j}$, magnetic flux density $\bm{B}$, magnetic field strength $\bm{H}$, hydrodynamic pressure $p$ and electric field strength $\bm{E}$, subject to body force field $\bm{f}$, a velocity initial condition $\bm{u}^{0}:=\bm{u}(\bm{x},0)$ and a divergence-free magnetic initial condition $\bm{B}^{0}:=\bm{B}(\bm{x},0)$, and proper boundary conditions [2], in a space-time domain, $\Omega\times(0,T]$. The material parameters involved are the fluid density $\rho$, the dynamic viscosity $\tilde{\mu}$, the electric conductivity $\sigma$, and the magnetic permeability $\mu$.
The incompressible MHD equations, (1), are a coupled system of the incompressible Navier-Stokes equations and the Maxwell’s equations. The two evolution equations, (1a) and (1c), stand for the conservation of momentum and the Faraday’s law of induction, respectively. Ohm’s law and Ampere’s law are (1d) and (1e), respectively. (1f) sets up a constitutive relation between $\bm{B}$ and $\bm{H}$. The divergence-free constraint on $\bm{u}$, (1b), represents mass conservation of the constant mass density flow. It is seen that, as $\bm{B}^{0}:=\bm{B}(\bm{x},0)$ is divergence free, (1c) ensures
$\nabla\cdot\bm{B}=0\quad\mathrm{in}\ \Omega\times(0,T],$
which is Gauss’s law for magnetism. And, from (1e), we know $\nabla\cdot\bm{j}=0$ which stands for conservation of charge. The coupling between the flow field and the electromagnetic field is through the Lorentz force, $\bm{j}\times\bm{B}$ in (1a), and the electromotive force, $\bm{u}\times\bm{B}$ in (1d).
Let $L$, $U$, and $B$ denote characteristic quantities of length, velocity, and magnetic flux density, respectively. We scale the space-time domain and variables by
$\bm{x}\leftarrow\dfrac{\bm{x}}{L},\quad t\leftarrow\dfrac{Ut}{L},$
$\bm{u}\leftarrow\dfrac{\bm{u}}{U},\quad p\leftarrow\dfrac{p}{\rho U^{2}},\quad\bm{f}\leftarrow\dfrac{L\bm{f}}{U^{2}},$
$\bm{B}\leftarrow\dfrac{\bm{B}}{B},\quad\bm{H}\leftarrow\dfrac{\mu\bm{H}}{B},\quad\bm{E}\leftarrow\dfrac{\bm{E}}{UB},\quad\bm{j}\leftarrow\dfrac{\mu L\bm{j}}{B}.$
And the induced scaled operators are
$D\leftarrow LD,\ D\in\left\{\nabla,\nabla\times,\nabla\cdot\right\},$
$\Delta\leftarrow L^{2}\Delta,\quad\partial_{t}\leftarrow\dfrac{L}{U}\partial_{t},\quad\bm{u}\cdot\nabla\leftarrow\dfrac{L}{U}\left(\bm{u}\cdot\nabla\right).$
Note that we do not introduce new notations for the scaled objects and, after this point, they always refer to the scaled dimensionless ones.
Now (1) can be written in a dimensionless form, i.e.,
(2a) $\partial_{t}\bm{u}+\bm{\omega}\times\bm{u}+\mathrm{R}_{f}^{-1}\nabla\times\bm{\omega}-\mathsf{c}\left(\nabla\times\bm{H}\right)\times\bm{H}+\nabla P$ $=\bm{f},$ in $\qquad\text{in }\Omega\times(0,T],$
(2b) $\bm{\omega}-\nabla\times\bm{u}$ $=\bm{0},$ in $\qquad\text{in }\Omega\times(0,T],$
(2c) $\nabla\cdot\bm{u}$ $=0,$ in $\qquad\text{in }\Omega\times(0,T],$
(2d) $\partial_{t}\bm{H}+\mathrm{R}_{m}^{-1}\Delta\bm{H}-\nabla\times\left(\bm{u}\times\bm{H}\right)$ $=\bm{0},$ in $\qquad\text{in }\Omega\times(0,T],$
[…]
(3a) $\partial_{t}\bm{B}+\nabla\times\bm{E}$ $=\bm{0},$
(3b) $\mathrm{R}_{m}^{-1}\bm{j}-\left(\bm{E}+\bm{u}\times\bm{B}\right)$ $=\bm{0},$
(3c) $\bm{j}-\nabla\times\bm{H}$ $=\bm{0},$
(3d) $\bm{B}$ $=\bm{H},$
have been written into (2d), which is the magnetic induction (or Helmholtz) equation. [^1] Note that, after non-dimensionalization, we get $\bm{B}=\bm{H}$, i.e. (3d). This is the case in vector calculus. In differential forms, they appear as a pair of differently oriented dual forms connected by the Hodge $\star$ operator, i.e., $\bm{B}=\star\bm{H}$. The three dimensionless parameters are the fluid Reynolds number $\mathrm{R}_{f}=\dfrac{\rho UL}{\tilde{\mu}}=\dfrac{UL}{\nu}$ (with $\nu=\dfrac{\tilde{\mu}}{\rho}$ being the kinematic viscosity), the magnetic Reynolds number $\mathrm{R}_{m}=\mu\sigma UL$, and the coupling number $\mathsf{c}=\mathrm{A}_{l}^{-2}$ ($\mathrm{A}_{l}=\dfrac{U\sqrt{\rho\mu}}{B}=\dfrac{U}{U_{\mathrm{A}}}$ is the Alfvén number and $U_{\mathrm{A}}=\dfrac{B}{\sqrt{\rho\mu}}$ is the Alfvén speed). Sometimes, the coupling term in (2a), $\mathsf{c}\left(\nabla\times\bm{H}\right)\times\bm{H}$, is written as $\mathsf{s}\left(\bm{E}+\bm{u}\times\bm{H}\right)\times\bm{H}$ where the coupling number becomes $\mathsf{s}=\mathsf{c}\mathrm{R}_{m}=\dfrac{\sigma LB^{2}}{\rho U}$. For example, see [3, (2.2a)].
To close the problem, (2) is complemented with initial conditions $\bm{u}^{0},\bm{H}^{0}$ and boundary conditions. In this work, we consider the following general boundary condition configuration:
  • 1.
    For the fluid part, there are two pairs of boundary conditions,
    (4) {$&P=\widehat{P}\quad&&\text{on }\Gamma_{\widehat{P}}\times\left(0,T\right]\\ &\bm{u}\cdot\bm{n}=\widehat{u}\quad&&\text{on }\Gamma_{\widehat{u}}\times\left(0,T\right]\end{aligned}\right.\quad\text{and}\quad\left\{\begin{aligned} &\bm{u}\times\bm{n}=\widehat{\bm{u}}\quad&&\text{on }\Gamma_{\widehat{\bm{u}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{\omega}\times\bm{n}\right)=\widehat{\bm{\omega}}\quad&&\text{on }\Gamma_{\widehat{\bm{\omega}}}\times\left(0,T\right]\end{aligned}\right.,$
    where $\bm{n}$ is the outward unit normal vector, $\left\{\Gamma_{\widehat{P}},\Gamma_{\widehat{u}}\right\}$ and $\left\{\Gamma_{\widehat{\bm{u}}},\Gamma_{\widehat{\bm{\omega}}}\right\}$ are two partitions of the boundary, such that $\overline{\Gamma}_{\widehat{P}}\cup\overline{\Gamma}_{\widehat{u}}=\partial\Omega$, $\Gamma_{\widehat{P}}\cap\Gamma_{\widehat{u}}=\emptyset$ and $\overline{\Gamma}_{\widehat{\bm{u}}}\cup\overline{\Gamma}_{\widehat{\bm{\omega}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{u}}}\cap\Gamma_{\widehat{\bm{\omega}}}=\emptyset$.
  • 2.
    For the Maxwell part, boundary conditions are given as
    (5) $&\bm{E}\times\bm{n}=\widehat{\bm{E}}\quad&&\text{on }\Gamma_{\widehat{\bm{E}}}\times\left(0,T\right]\\ &\bm{n}\times\left(\bm{H}\times\bm{n}\right)=\widehat{\bm{H}}\quad&&\text{on }\Gamma_{\widehat{\bm{H}}}\times\left(0,T\right]\end{aligned}\right.,$
    where $\left\{\Gamma_{\widehat{\bm{E}}},\Gamma_{\widehat{\bm{H}}}\right\}$ is a partition of the boundary, satisfying $\overline{\Gamma}_{\widehat{\bm{E}}}\cup\overline{\Gamma}_{\widehat{\bm{H}}}=\partial\Omega$, $\Gamma_{\widehat{\bm{E}}}\cap\Gamma_{\widehat{\bm{H}}}=\emptyset$. Note that the boundary condition $\bm{E}\times\bm{n}$ implies the magnetic flux boundary condition $\bm{H}\cdot\bm{n}$ (or $\bm{B}\cdot\bm{n}$). This is because
[…]
    on $\partial\Omega\times\left(0,T\right]$. Also see [4, Section 2.3].
For example, the boundary conditions considered in the work of Hu et al., see (7) of [5], are a particular case of this general configuration, i.e. the case when $\Gamma_{\widehat{u}}=\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset$, $\widehat{P}=0$ and $\widehat{\bm{u}}=\widehat{\bm{E}}=\bm{0}$.
It is generally acknowledged that structure-preservation, especially the preservation (either strongly of weakly) of $\nabla\cdot\bm{B}=0$, is an essential property of numerical methods for MHD [6]. Many of these structure-preserving methods can be classified into, for example, potential-based methods [7], divergence-cleaning methods [8, 9], constrained transport methods [10, 11] and Lagrange multiplier (or augmented) methods [12, 13, 14]. In the current paper neither of those techniques will be applied, but it will be shown that the absence of magnetic monopoles is weakly enforced through the decoupled variational formulation. Recently, methods that aim at preserving multiple structures in the mixed finite element setting become increasingly popular. See, for example, the work of Hu et al. for incompressible MHD that preserves cross- and magnetic-helicity, energy and Gauss law of magnetism [5], the work of Gawlik and Gay-Balmaz for incompressible MHD of a variable fluid density that preserves energy, cross-helicity (when the fluid density is constant) and magnetic-helicity, mass, total squared density, pointwise incompressibility, and Gauss’s law for magnetism [15], the work of Laakmann et al. [16] and references therein. These methods preserve several physical quantities of interest, but usually are less computationally efficient because of their large (nonlinear) discrete systems to be solved. A compromise is to use a decoupled scheme which breaks the discrete system into several smaller ones, for example see [3, 12, 13, 17, 18]. These methods in the mixed finite element setting commonly deal with a particular configuration of boundary conditions.
In this paper, we introduce a structure-preserving mixed finite element discretization for the incompressible MHD under a general boundary condition configuration. The method is decoupled in the sense that the two evolution equations are temporally discretized at different time sequences such that they can be solved individually.
The outline of the rest of the paper is as follows. Relevant Hilbert spaces and the de Rham Hilbert complex are introduced in Section 2 followed by the proposed semi-discrete formulation and its conservation properties in Section 3. The temporal discretization that leads to the decoupled scheme is explained in Section 4. Numerical tests are presented in Section 5. Finally, in Section 6 conclusions are drawn.
2 Relevant Hilbert spaces and the de Rham complex
The function spaces used in this work are similar to those used in [19] where only the two-dimensional setting is discussed. For the sake of completeness, we give a brief introduction to the three-dimensional spaces as a complement to [19, Section 2.1]. We first introduce the relevant spaces and the de Rham complex in the infinite dimensional setting, and then introduce their finite dimensional counterparts. For an extensive introduction of function spaces, see [20] or any other textbook on functional analysis.
2.1 The infinite dimensional setting
Let $L^{2}\left(\Omega\right)$ denote the space of square-integrable functions on $\Omega$,
$L^{2}(\Omega):=\left\{\varphi\left|\left\langle{\varphi},{\varphi}\right\rangle_{\Omega}:=\int_{\Omega}\varphi\cdot\varphi\ \mathrm{d}\Omega<\infty\right.\right\}.$
For $\Omega\subset\mathbb{R}^{3}$, we also consider the following Hilbert spaces,
$H^{1}\left(\Omega\right)$ :=$:=\left\{\psi\left|\psi\in L^{2}(\Omega),\nabla\psi\in\left[L^{2}\left(\Omega\right)\right]^{3}\right.\right\},$
$H\left(\mathrm{curl};\Omega\right)$ :=$:=\left\{\bm{\sigma}\left|\bm{\sigma}\in\left[L^{2}\left(\Omega\right)\right]^{3},\nabla\times\bm{\sigma}\in\left[L^{2}\left(\Omega\right)\right]^{3}\right.\right\},$
$H\left(\mathrm{div};\Omega\right)$ :=$:=\left\{\bm{\phi}\left|\bm{\phi}\in\left[L^{2}\left(\Omega\right)\right]^{3},\nabla\cdot\bm{\phi}\in L^{2}\left(\Omega\right)\right.\right\}.$
They form the well known de Rham Hilbert complex [21, 22],
(7) $\mathbb{R}\hookrightarrow H^{1}\left(\Omega\right)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}H(\mathrm{curl};\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}H(\mathrm{div};\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}L^{2}(\Omega)\to 0.$
[…]
The trace operator, $\mathcal{T}$, restricts an element of Hilbert spaces to a boundary section $\Gamma\subseteq\partial\Omega$. For $\psi\in H^{1}\left(\Omega\right)$ and $\bm{\phi}\in H\left(\mathrm{div};\Omega\right)$, we define its action as
$\mathcal{T}\psi=\left.\psi\right|_{\Gamma},\quad\mathcal{T}\bm{\phi}=\left.\bm{\phi}\cdot\bm{n}\right|_{\Gamma}.$
And we distinguish between $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ for elements of $H\left(\mathrm{curl};\Omega\right)$, i.e., for $\bm{\sigma},\bm{\eta}\in H\left(\mathrm{curl};\Omega\right)$,
(8) $\mathcal{T}\bm{\sigma}=\left.\bm{\sigma}\times\bm{n}\right|_{\Gamma},\quad\mathcal{T}_{\parallel}\bm{\eta}=\left.\bm{n}\times\left(\bm{\eta}\times\bm{n}\right)\right|_{\Gamma}.$
Using the above, we define the trace spaces on $\Gamma$ as
$H^{1/2}\left(\Omega,\Gamma\right):=\left\{\mathcal{T}\psi\left|\psi\in H^{1}(\Omega)\right.\right\},$
$\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right):=\left\{\mathcal{T}\bm{\sigma}\left|\bm{\sigma}\in H(\mathrm{curl};\Omega)\right.\right\},$
$\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right):=\left\{\mathcal{T}_{\parallel}\bm{\eta}\left|\bm{\eta}\in H(\mathrm{curl};\Omega)\right.\right\},$
$\mathcal{T}H\left(\mathrm{div};\Omega,\Gamma\right):=\left\{\mathcal{T}\bm{\phi}\left|\bm{\phi}\in H(\mathrm{div};\Omega)\right.\right\}.$
Note that, at the continuous level, although $\mathcal{T}$ and $\mathcal{T}_{\parallel}$ are different, see (8), trace spaces $\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma\right)$ and $\mathcal{T}_{\parallel}H\left(\mathrm{curl};\Omega,\Gamma\right)$ refer to the same tangential space.
2.2 The finite dimensional setting
Assume that we are given finite dimensional spaces $G(\Omega)\subset H^{1}(\Omega)$, $C(\Omega)\subset H(\mathrm{curl};\Omega)$, $D(\Omega)\subset H(\mathrm{div};\Omega)$, $S(\Omega)\subset L^{2}(\Omega)$, which form a discretization of the continuous de Rham Hilbert complex,
(9) $\mathbb{R}\hookrightarrow G(\Omega)\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}C(\Omega)\stackrel{{\scriptstyle\nabla\times}}{{\longrightarrow}}D(\Omega)\stackrel{{\scriptstyle\nabla\cdot}}{{\longrightarrow}}S(\Omega)\to 0.$
In particular, this means that if the complex in (7) is exact, then so is the one in (9). We will also consider the following subspaces with boundary conditions, where $\Gamma\subset\partial\Omega$,
$G_{\widehat{\psi}}(\Omega,\Gamma):=\left\{\psi_{h}\left|\psi_{h}\in G(\Omega),\ \mathcal{T}\psi_{h}=\widehat{\psi}\ \text{on}\ \Gamma\right.\right\},$
$C_{\bm{\widehat{\sigma}}}(\Omega,\Gamma):=\left\{\bm{\sigma}_{h}\left|\bm{\sigma}_{h}\in C(\Omega),\ \mathcal{T}\bm{\sigma}_{h}=\bm{\widehat{\sigma}}\ \text{on}\ \Gamma\right.\right\},$
$C^{\parallel}_{\bm{\widehat{\eta}}}(\Omega,\Gamma):=\left\{\bm{\eta}_{h}\left|\bm{\eta}_{h}\in C(\Omega),\ \mathcal{T}_{\parallel}\bm{\eta}_{h}=\bm{\widehat{\eta}}\ \text{on}\ \Gamma\right.\right\},$
$D_{\widehat{\phi}}(\Omega,\Gamma):=\left\{\bm{\phi}_{h}\left|\bm{\phi}_{h}\in D(\Omega),\ \mathcal{T}\bm{\phi}_{h}=\widehat{\phi}\ \text{on}\ \Gamma\right.\right\}.$
For the corresponding homogeneous boundary conditions we have
$G_{0}(\Omega,\Gamma):=\left\{\psi_{h}\left|\psi_{h}\in G(\Omega),\ \mathcal{T}\psi_{h}=0\ \text{on}\ \Gamma\right.\right\},$
$C_{\bm{0}}(\Omega,\Gamma):=\left\{\bm{\sigma}_{h}\left|\bm{\sigma}_{h}\in C(\Omega),\ \mathcal{T}\bm{\sigma}_{h}=\bm{0}\ \text{on}\ \Gamma\right.\right\},$
$C^{\parallel}_{\bm{0}}(\Omega,\Gamma):=\left\{\bm{\eta}_{h}\left|\bm{\eta}_{h}\in C(\Omega),\ \mathcal{T}_{\parallel}\bm{\eta}_{h}=\bm{0}\ \text{on}\ \Gamma\right.\right\},$
$D_{0}(\Omega,\Gamma):=\left\{\bm{\phi}_{h}\left|\bm{\phi}_{h}\in D(\Omega),\ \mathcal{T}\bm{\phi}_{h}=0\ \text{on}\ \Gamma\right.\right\}.$
Moreover, the finite dimensional trace spaces will be denoted by $\mathcal{T}G(\Omega,\Gamma)$, $\mathcal{T}C(\Omega,\Gamma)$, $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$, $\mathcal{T}D(\Omega,\Gamma)$. Unlike their continuous counterparts, finite dimensional trace spaces $\mathcal{T}C(\Omega,\Gamma)$ and $\mathcal{T}_{\parallel}C(\Omega,\Gamma)$ usually are different.
We define a corresponding trilinear form,
(10) $a\left(\bm{\alpha}_{h},\bm{\beta}_{h},\bm{\gamma}_{h}\right):=\left\langle{\bm{\alpha}_{h}\times\bm{\beta}_{h}},{\bm{\gamma}_{h}}\right\rangle_{\Omega},\quad\bm{\alpha}_{h},\bm{\beta}_{h},\bm{\gamma}_{h}\in\left\{C(\Omega),D(\Omega)\right\},$
which is skew-symmetric with respect to any two of the three entries. Note that, in the infinite dimensional setting, the $L^{2}$-boundedness of $\left\langle{\bm{\alpha}\times\bm{\beta}},{\bm{\gamma}}\right\rangle_{\Omega}$ is not implied even if $\bm{\alpha}$, $\bm{\beta}$ and $\bm{\gamma}$ are $L^{2}$ vector fields. However, in the finite dimensional setting where the spaces $C(\Omega)$ and $D(\Omega)$ consist of piecewise-polynomial finite element vector fields, it is in fact $L^{2}$-bounded.
3 Spatially-discrete formulation and its conservation properties
In this section, we present a spatially-discrete weak formulation of (2) and analyze its conservation properties. As a remark, for neatness, we only discuss the three-dimensional formulation and its temporal discretization in this paper. Obtaining their two-dimensional versions is straightforward.
3.1 Spatially-discrete formulation
To increase the notational clarity, we omit the part for the temporal domain. The proposed semi-discrete weak formulation of (2) in $\mathbb{R}^{3}$ is written as: Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{H}^{0}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)$, seek $\left(\bm{u}_{h},\bm{\omega}_{h},P_{h},\bm{H}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C^{\parallel}_{\widehat{\bm{\omega}}}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)\times C^{\parallel}_{\widehat{\bm{H}}}(\Omega,\Gamma_{\widehat{\bm{H}}})$, such that $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h},\bm{b}_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}_{\perp}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)\times C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$,
(11a) $\left\langle\partial_{t}\bm{u}_{h},\bm{v}_{h}\right\rangle_{\Omega}+a\left(\bm{\omega}_{h},\bm{u}_{h},\bm{v}_{h}\right)+\mathrm{R}_{f}^{-1}\left\langle{\nabla\times\bm{\omega}_{h}},{\bm{v}_{h}}\right\rangle_{\Omega}\qquad$
$-\mathsf{c}\ a\left(\nabla\times\bm{H}_{h},\bm{H}_{h},\bm{v}_{h}\right)-\left\langle{P_{h}},{\nabla\cdot\bm{v}_{h}}\right\rangle_{\Omega}$ $=\left\langle{\bm{f}},{\bm{v}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}},{\mathcal{T}\bm{v}_{h}}\right\rangle_{\Gamma_{\widehat{P}}},$
(11b) $-\left\langle{\bm{u}_{h}},{\nabla\times\bm{w}_{h}}\right\rangle_{\Omega}+\left\langle{\bm{\omega}_{h}},{\bm{w}_{h}}\right\rangle_{\Omega}$ $=-\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{w}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{u}}}},$
(11c) $\left\langle{\nabla\cdot\bm{u}_{h}},{q_{h}}\right\rangle_{\Omega}$ $=0,$
(11d) $\left\langle{\partial_{t}\bm{H}_{h}},{\bm{b}_{h}}\right\rangle_{\Omega}+\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\bm{H}_{h}},{\nabla\times\bm{b}_{h}}\right\rangle_{\Omega}-\ a\left(\bm{u}_{h},\bm{H}_{h},\nabla\times\bm{b}_{h}\right)$ $=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\bm{b}_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}.$
Note that this assumes that the supplied boundary conditions are exactly representable as members of the chosen finite dimensional spaces; more generally, one can always project the boundary conditions into the appropriate finite dimensional trace spaces. See (4) and (5) for the general boundary setting.
3.2 Conservation properties
In this section, we analyze some properties of the formulation (11). In particular, we show that this semi-discrete formulation preserves conservation of mass and conservation of charge strongly, preserves Gauss’ law of magnetism weakly, and preserves the correct energy dissipation rate such that, when no body forces are applied and there is no net flux of energy through the boundary, it conserves the total energy in the ideal limit.
3.2.1 Conservation of mass
[…]
3.2.2 Conservation of charge
Since we have selected $\bm{H}_{h}\in C(\Omega)$, from (9), we know that we can always find $\bm{j}_{h}=\nabla\times\bm{H}_{h}\in D(\Omega)$ which is pointwise divergence-free; conservation of charge is preserved by the formulation (11).
3.2.3 Weak conservation of Gauss’s law for magnetism
Since (11d) is valid for all $\bm{b}_{h}\in C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$, let us choose $\bm{b}_{h}=\nabla q_{h}$ for an arbitrary $q_{h}\in G_{0}(\Omega,\Gamma_{\widehat{\bm{H}}})$. Then, we have
(12) $\left\langle{\partial_{t}\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}+\mathrm{R}_{m}^{-1}\left\langle{\nabla\times\bm{H}_{h}},{\nabla\times\nabla q_{h}}\right\rangle_{\Omega}-\ a\left(\bm{u}_{h},\bm{H}_{h},\nabla\times\nabla q_{h}\right)=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\nabla q_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\;.$
The second and third terms on the left hand side are zero because we strongly satisfy ×$\nabla\times\nabla q_{h}=0$ since our discrete function spaces constitute a finite dimensional de Rham complex. We are left with the following equation,
(13) $\left\langle{\partial_{t}\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\nabla q_{h}}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\,.$
Using the fact that $\nabla q_{h}\in C_{\bm{0}}^{\parallel}\left(\Omega,\Gamma_{\widehat{\bm{H}}}\right)$, we have
(14) $\left\langle{\partial_{t}\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}=\left\langle\widehat{\bm{E}},\mathcal{T}_{\parallel}\nabla q_{h}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}$ $=\left\langle\widehat{\bm{E}},\mathcal{T}_{\parallel}\nabla q_{h}\right\rangle_{\partial\Omega}$
$=-\left\langle\left(\nabla\cdot\right)_{\partial\Omega}\widehat{\bm{E}},\mathcal{T}q_{h}\right\rangle_{\partial\Omega}=-\left\langle\left(\nabla\cdot\right)_{\Gamma_{\widehat{\bm{E}}}}\widehat{\bm{E}},\mathcal{T}q_{h}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\;.$
Finally, using the relation between $\bm{H}\cdot\bm{n}$ and $\widehat{\bm{E}}$ from (6), we can rewrite the rightmost term to obtain the following equation,
(15) $\dfrac{\mathrm{d}}{\mathrm{d}t}\left[\left\langle{\bm{H}_{h}},{\nabla q_{h}}\right\rangle_{\Omega}-\left\langle\bm{H}\cdot\bm{n},\mathcal{T}q_{h}\right\rangle_{\Gamma_{\widehat{\bm{E}}}}\right]=0\;.$
The expression inside the square brackets is precisely the weak divergence-free constraint on $\bm{H}_{h}$. The above equation thus states that the time-derivative of this constraint on $\bm{H}_{h}$ is zero, and implies that $\bm{H}_{h}$ will remain weakly divergence-free for all $t$ if the initial condition $\bm{H}^{0}_{h}$ is weakly divergence-free.
3.2.4 Energy Conservation
The (semi-)discrete (total) energy is defined as
(16) $\mathcal{E}_{h}:=\mathcal{K}_{h}+\mathcal{M}_{h}$
whose $\mathcal{K}_{h}:=\dfrac{1}{2}\left\langle{\bm{u}_{h}},{\bm{u}_{h}}\right\rangle_{\Omega}$ is the kinetic energy of the fluid and $\mathcal{M}_{h}:=\dfrac{\mathsf{c}}{2}\left\langle{\bm{H}_{h}},{\bm{H}_{h}}\right\rangle_{\Omega}$ is the magnetic energy. Suppose
(17) $\widehat{u}=0\;,\quad\Gamma_{\widehat{\bm{u}}}=\Gamma_{\widehat{\bm{E}}}=\partial\Omega\;,\quad\Gamma_{\widehat{\bm{\omega}}}=\Gamma_{\widehat{\bm{H}}}=\emptyset\;.$
By replacing $\bm{v}_{h}$ in (11a) with $\bm{u}_{h}$, we can obtain
(18) $\partial_{t}\mathcal{K}_{h}+\mathrm{R}_{f}^{-1}\left\langle{\nabla\times\bm{\omega}_{h}},{\bm{u}_{h}}\right\rangle_{\Omega}-\mathsf{c}\ a\left(\bm{j}_{h},\bm{H}_{h},\bm{u}_{h}\right)=\left\langle{\bm{f}},{\bm{u}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}},{\mathcal{T}\bm{u}_{h}}\right\rangle_{\Gamma_{\widehat{P}}},$
where $\bm{j}_{h}=\nabla\times\bm{H}_{h}\in D(\Omega)$, and the fluid advection term $a\left(\bm{\omega}_{h},\bm{u}_{h},\bm{u}_{h}\right)$ and the total pressure term $\left\langle{P_{h}},{\nabla\cdot\bm{u}_{h}}\right\rangle_{\Omega}$ have vanished because of the skew-symmetry of the trilinear (10) and $\nabla\cdot\bm{u}_{h}=0$, respectively. Meanwhile, because (11b) holds $\forall\bm{w}_{h}\in C(\Omega)$ and $\bm{\omega}_{h}\in C(\Omega)$, we know that
$-\left\langle{\bm{u}_{h}},{\nabla\times\bm{\omega}_{h}}\right\rangle_{\Omega}+\left\langle{\bm{\omega}_{h}},{\bm{\omega}_{h}}\right\rangle_{\Omega}=-\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{\omega}_{h}}\right\rangle_{\partial\Omega}.$
As a results, (18) can be expressed as
(19) $\partial_{t}\mathcal{K}_{h}+\mathrm{R}_{f}^{-1}\left\langle{\bm{\omega}_{h}},{\bm{\omega}_{h}}\right\rangle_{\Omega}-\mathsf{c}\ a\left(\bm{j}_{h},\bm{H}_{h},\bm{u}_{h}\right)=\left\langle{\bm{f}},{\bm{u}_{h}}\right\rangle_{\Omega}-\left\langle{\widehat{P}},{\mathcal{T}\bm{u}_{h}}\right\rangle_{\Gamma_{\widehat{P}}}+\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{\omega}_{h}}\right\rangle_{\partial\Omega}.$
Similarly, replacing $\bm{b}_{h}$ in (11d) with $\bm{H}_{h}$ leads to
(20) $\dfrac{1}{\mathsf{c}}\partial_{t}\mathcal{M}_{h}+\mathrm{R}_{m}^{-1}\left\langle{\bm{j}_{h}},{\bm{j}_{h}}\right\rangle_{\Omega}-a\left(\bm{u}_{h},\bm{H}_{h},\bm{j}_{h}\right)=\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\bm{H}_{h}}\right\rangle_{\partial\Omega}.$
With (19) and (20), we can finally derive the rate of change of the discrete energy,
(21) $\partial_{t}\mathcal{E}_{h}$ $=\partial_{t}\mathcal{K}_{h}+\partial_{t}\mathcal{M}_{h}$
$=\underbrace{\left\langle{\bm{f}},{\bm{u}_{h}}\right\rangle_{\Omega}}_{\text{(i)}}\ \underbrace{-\mathrm{R}_{f}^{-1}\mathcal{S}_{h}}_{\text{(ii)}}\ \underbrace{-\mathsf{c}\mathrm{R}_{m}^{-1}\mathcal{J}_{h}}_{\text{(iii)}}\underbrace{-\left\langle{\widehat{P}},{\mathcal{T}\bm{u}_{h}}\right\rangle_{\Gamma_{\widehat{P}}}}_{\text{(iv)}}+\underbrace{\left\langle{\widehat{\bm{u}}},{\mathcal{T}_{\parallel}\bm{\omega}_{h}}\right\rangle_{\partial\Omega}}_{\text{(v)}}+\underbrace{\mathsf{c}\left\langle{\widehat{\bm{E}}},{\mathcal{T}_{\parallel}\bm{H}_{h}}\right\rangle_{\partial\Omega}}_{\text{(vi)}},$
where $\mathcal{S}_{h}:=\left\langle{\bm{\omega}_{h}},{\bm{\omega}_{h}}\right\rangle_{\Omega}$, $\mathcal{J}_{h}:=\left\langle{\bm{j}_{h}},{\bm{j}_{h}}\right\rangle_{\Omega}$. The two nonlinear terms have canceled each other due to the skew-symmetry of the trilinear form (10); the exchange of energy between $\mathcal{K}_{h}$ and $\mathcal{M}_{h}$ is exactly captured.
In (21), (i) represents amount of mechanical work done on the system per unit time, (ii) and (iii) are dissipation rates due to viscosity and electric resistance, respectively, and terms (iv) - (vi) are the net flux of energy through the domain boundary.
Clearly, if $\bm{f}=\bm{0}$, when the MHD flow is ideal (i.e. $\mathrm{R}_{f}=\mathrm{R}_{m}=\infty$) and there is no net flux of energy through the domain boundary, formulation (11) preserves energy, i.e.,
$\partial_{t}\mathcal{E}_{h}=\partial_{t}\mathcal{K}_{h}+\partial_{t}\mathcal{M}_{h}=0.$
4 Fully-discrete weak formulation
If a conservative temporal discrete scheme, for example, the Crank-Nicolson scheme, is applied to the semi-discrete formulation (11) directly, we can obtain a fully-discrete formulation of same conservation and dissipation properties (except that the time derivative of energy is now expressed in a temporally discrete format). The proof is straightforward; it mimics the analyses in Section 3.2. Therefore, the rest of this paper is devoted to a decoupled formulation with computational efficiency in mind.
4.1 Decoupled temporal discretization
Consider a set of time instants,
$\left\{t^{0},t^{\frac{1}{2}},t^{1},t^{1+\frac{1}{2}},t^{2},\cdots\right\},$
where $t^{0}=0$, $\Delta t=t^{k}-t^{k-1}>0$, $k\in\left\{1,2,3,\cdots\right\}$, is a constant time interval, and $t^{k-\frac{1}{2}}=\dfrac{t^{k-1}+t^{k}}{2}$. We use a superscript to denote the evaluation of a variable at a particular time instant. For example, $\bm{u}^{k}_{h}:=\bm{u}_{h}\left(\bm{x},t^{k}\right)$.
A fully discrete, temporally decoupled formulation of (11) (hereinafter referred to as the decoupled formulation) is expressed as follows. Given $\bm{f}\in\left[L^{2}(\Omega)\right]^{3}$, natural boundary conditions $\widehat{P}\in H^{1/2}\left(\Omega;\Gamma_{\widehat{P}}\right),\ \widehat{\bm{u}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{u}}}\right),\ \widehat{\bm{E}}\in\mathcal{T}H\left(\mathrm{curl};\Omega,\Gamma_{\widehat{\bm{E}}}\right)$, and initial conditions $\left(\bm{u}_{h}^{0},\bm{\omega}_{h}^{0},\bm{H}^{\frac{1}{2}}_{h}\right)\in D\left(\Omega\right)\times C\left(\Omega\right)\times C\left(\Omega\right)$, for time step index $k=1,2,3,\cdots$ successively,
(step 1) seek $\left(\bm{u}^{k}_{h},\bm{\omega}^{k}_{h},P^{k-\frac{1}{2}}_{h}\right)\in D_{\widehat{u}}(\Omega,\Gamma_{\widehat{u}})\times C_{\widehat{\bm{\omega}}}^{\parallel}(\Omega,\Gamma_{\widehat{\bm{\omega}}})\times S(\Omega)$, such that, $\forall\left(\bm{v}_{h},\bm{w}_{h},q_{h}\right)\in D_{0}\left(\Omega,\Gamma_{\widehat{u}}\right)\times C^{\parallel}_{\bm{0}}\left(\Omega,\Gamma_{\widehat{\bm{\omega}}}\right)\times S\left(\Omega\right)$,
(22a) $\left\langle\dfrac{\bm{u}^{k}_{h}-\bm{u}^{k-1}_{h}}{\Delta t},\bm{v}_{h}\right\rangle_{\Omega}+a\left(\dfrac{\bm{\omega}^{k-1}_{h}+\bm{\omega}^{k}_{h}}{2},\dfrac{\bm{u}^{k-1}_{h}+\bm{u}^{k}_{h}}{2},\bm{v}_{h}\right)$ +Rf1